OpenStudy (anonymous):

I need a little help with: Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = −2

OpenStudy (anonymous):

my attempt 2x + y = 1 4x + 2y = −2 2x + y = 1 -2x -2x y = 1 -2x 4x + 2y = −2 4x + 2(1-2x) = −2 4x + 2 -4x=-2 x +2=-2 -2 -2 ____________ x=-4 2(-4) + y = 1 -8+y=1 +8 +8 _____________ y = 7 x=-4 ______________ checking 2(-4)+ 7 = 1 -8+7=1 -1=1 4x + 2y = −2 4(-4) + 2(7) = −2 -16 + 14 =-2 -2=-2

OpenStudy (anonymous):

This system of equations has no solution and is graphically a pair of parallel lines.

OpenStudy (anonymous):

Your attempt was going well up until your 9th line. At that point, the x terms completely disappear.

OpenStudy (anonymous):

Another way to see that you have parallel lines is to put the equations into slope-intercept form. There you would see that you have 2 equations of the form: y=mx+b where the "m" values are the same but the "b" s are different.

OpenStudy (anonymous):

inst 4+-4 =0 so nothing goes by the x?

OpenStudy (anonymous):

y = -2x + 1 and y = -2x - 1

OpenStudy (anonymous):

2x + y = 1 => y = 1-2x 4x + 2y = -2 4x + 2 (1 -2x) = -2 4x + 2 - 4x = -2 2= -2

OpenStudy (anonymous):



OpenStudy (anonymous):

Actually, your 9th line is not: x +2=-2 It instead becomes the incongruous: 2 = -2 which cannot be, which shows parallel lines. It is an "inconsistent" system.

OpenStudy (anonymous):

Hopefully, your homework assignment contains at least a couple of more sets of equations, some pairs of which that can be solved, but this set definitely cannot be solved.

OpenStudy (anonymous):

You see, in your 8th line, you have 4x and -4x on the left side. They cancel and you are left with no "x" terms.

OpenStudy (anonymous):

So, your 9th line is not: x +2=-2 Instead, it is : 2 = -2 And that's where you can stop. You don't have to go further.

OpenStudy (anonymous):

When you are trying to solve a pair of x,y equations, you are going to get one of 3 situations to occur: 1) they intersect in one point (independent system) 2) they are parallel (inconsistent system) 3) they are actually the same line (dependent system) Here, you have #2, an inconsistent or parallel line system.

OpenStudy (anonymous):

Ah I see thank you. the assessment has the option of no solution So if 2= -2 I makes it parallel.